A movie theater charges $ 9 for adults and $ 6 for seniors. On a particular day when 334 people paid an admission, the total receipts were $ 2551. Round your answers to the nearest integers.
How many who paid were adults?
How many who paid were seniors?

Answer by macston(5194)

A=number of Adults, S=number of Seniors
9A=revenue from Adults, 6S=revenue from Seniors
On a day when 334 paid Admission
T=Total paid admission=A + S =334 Eq1
R=Total Receipts=9A + 6S=2551 Eq2
Multiply both sides of Eq1 by 9:
(9)(A + S) = (9)334
9A + 9S=3006 Eq3 Subtract Eq2 from Eq3
(9A + 9S)-(9A +6S)=3006-2551
9A+9S-9A-6S=455
3S=455
S=152 ANSWER 1 Number of Paid Seniors=152
A+S=334
A+152=334
A+152-152=334-152 Subtract 152 from both sides
A=182 ANSWER 2 Number of Paid Adults=182
CHECK:
Total receipts=9A+6S=2551
9(182)+6(152)=2551
1638+912=2551
2550=2551 The difference here attributed to rounding